Match the probability with the outcome if two number cubes are rolled 100 times.

two 6s appeared 3 times

Answer by stanbon(48568)

If you are only interested in the theoretical versus the empirical
probability for two-sixes you have:
empirical = 3/1000= 0.00300
---------
expected number of double-sixes is
1000*(1/36) = 37.78
----------------------
Probability of double-sixes = 37.78/1000 = 0.0378
==========================
Cheers,
Stan H.